a ring and a disc of same mass roll without slipping along a horizontal surface with same velocity. if the kinetic energy of ring is 8J then that of disc is

ans:6J
1
Expert's answer
2019-04-02T10:57:36-0400

The ring and the disc which are rolling without slipping take part both the translational and rotational movements simultaneously, therefore their full kinetic energy is

"W=\frac{M{{v}^{2}}}{2}+\frac{J{{\omega }^{2}}}{2}"

where

"\frac{M{{v}^{2}}}{2}={{W}_{T}}"

is the Translational Kinetic Energy

"\frac{J{{\omega }^{2}}}{2}={{W}_{R}}"

is the Rotational Kinetic Energy, M is the mass, v is the velocity, J is the moment of inertia, ω is the angular velocity.

The moments of inertia of the ring and the disc with radii Rr and Rd are

"J=MR_{r}^{2},\,\,J=\frac{MR_{d}^{2}}{2}"

respectively.

Then the Kinetic Energy of the ring Wr is

"{{W}_{r}}=\frac{M{{v}^{2}}}{2}+\frac{MR_{r}^{2}\omega _{r}^{2}}{2}"

When rotating without slipping we have

"{{R}_{r}}{{\omega }_{r}}={{v}_{r}}=v"

Then


"{{W}_{r}}=\frac{M{{v}^{2}}}{2}+\frac{M{{v}^{2}}}{2}=M{{v}^{2}}=8J"

The Kinetic Energy of the disc Wd is

"{{W}_{r}}=\frac{M{{v}^{2}}}{2}+\frac{MR_{d}^{2}\omega _{d}^{2}}{4}=\frac{M{{v}^{2}}}{2}+\frac{M{{v}^{2}}}{4}=\frac{3M{{v}^{2}}}{4}"

Since

"M{{v}^{2}}=8J"

then


"{{W}_{r}}=\frac{3M{{v}^{2}}}{4}=\frac{3}{4}\cdot 8=6J"

Thus, the Kinetic Energy of disc is 6J.